Consider the following numerical example of the simple keynesian model with no government spending ,taxes or foreign sector (all figures in R millions):
C=100 + 0,9Y
I=50
Answer the following questions.
1-calculate the equilibrium level of output.
2-in equilibrium , what is the value of consumption spending ?use this number to verify that the sum of C and I in equilbrium equals the value for equilibrium output you obtained above.
3-what is the value of the multiplier in this economy?
1
Expert's answer
2019-08-13T08:56:26-0400

1. The equilibrium level of output is:

Y = C + I = 100 + 0.9Y + 50,

0.1Y = 150,

Y = 1500.

2.The value of consumption spending: C = 100+0.9*1500 = 1450

use this number to verify that the sum of C and I in equilbrium equals the value for equilibrium output you obtained above:

Y = C + I = 100 + 0.9Y + 50 = 100+0.9*1500+50 = 1500

3. The value of the marginal propensity to consume is: MPC = 0.9., then the value of the multiplier in this economy is:

1/(1-0.9) = 10